The cost in dollars of manufacturing w wing nuts is given by the expression 0.07w+13.3. find the cost when 375 wing nuts are made. what is the average cost to manufacture one wing nut?
Answer by checkley75(3666)

0.07W+13.3=0.07*375+13.3=26.25+13.3=39.55 COST OF 375 WING NUTS.
THUS THE COST PER WING NUT IS:
39.55/375=.105 FOR EACH WING NUT.